# Geometric-Aware Low-Light Image and Video Enhancement via Depth Guidance

### v1

@@ 3.3 Correlation-based Cross Attention for Fusion
$A_{l}=Softmax(W_{q}(f_{g,l})\times W_{k}(f_{g,l})^{T}\times\tau),$ (5)
where $W_{q}(f_{g,l})$ and $W_{k}(f_{g,l})$ has the shape of $C\times(H\times W)$, $T$ is the transpose operation, $\tau$ is the temperature value that is a hyper-parameter, $\times$ is the matrix multiplication operation, and $A_{l}$ has the shape of $C\times C$. The output of the self-attention is obtained as
$f^{\prime}_{g,l}=A_{l}\times W_{v}(f_{g,l}).$ (6)
@@
$f^{\prime\prime}_{g,l}=\hat{A}_{l}\times\hat{W}_{v}(\hat{f}_{d,l}).$
As both $f_{g,l}$ and $\hat{f}{d,l}$ are derived from the encoder process of $I_{d}$ and exhibit homogeneity, as discussed in Sec. 3.2, $f^{\prime}{g,l}$ and $f^{\prime\prime}{g,l}$ also exhibit small heterogeneity between them. To further minimize the disparity between $f^{\prime}{g,l}$ and $f^{\prime\prime}_{g,l}$, we introduce a method to model the correlation between these two sets of features and utilize this correlation as a weighting factor for the fusion process. The correlation is obtained as
$w_{l}=Sigmoid(\mathcal{O}(f^{\prime}_{g,l}\oplus f^{\prime\prime}_{g,l})),$ (8)
where $\oplus$ is the channel concatenation operation. The fusion output can be written as
$\bar{f}_{g,l}=w_{l}\cdot f^{\prime\prime}_{g,l}+f^{\prime}_{g,l}+f_{g,l},$ (9)
@@
3.4 Training Strategy
In Eq. 3, two objectives are considered: the restoration loss for enhancement and the depth supervision loss to obtain accurate depth priors. The restoration loss, denoted as $\mathcal{L}_{g}$, can employ the same loss terms as those used for the target model $\mathcal{G}$. On the other hand, the depth supervision loss is implemented as the $L_{2}$ distance between the predicted depth values and the ground truth depth information, as
$\mathcal{L}_{d}=\|\hat{d}_{d}-d_{n}\|_{2}.$ (11)
@@ 3.4 Training Strategy
$\mathcal{L}=\mathcal{L}_{g}+\lambda\mathcal{L}_{d},$ (12)
@@
where $\lambda$ is the loss weight, which is robust across different target scenarios. We guarantee that our code and models will be released upon the publication of this paper.
4 Experiment
4.1 Experimental Settings
Datasets. We assessed the proposed framework using various LLIE/LLVE datasets, which include SDSD-indoor [19], SDSD-outdoor [19], LOLv2 [30], SMID [3], SID [2], and DAVIS2017 [11]. SDSD: The videos in SDSD were captured in dynamic pairs using electromechanical equipment. Consequently, it is employed in both LLIE and LLVE tasks. LOLv2: LOLv2 is subdivided into LOLv2-real and LOLv2-synthetic. LOLv2-real comprises real-world low-/normal-light image pairs for training and testing. LOLv2-synthetic was generated by analyzing the illumination distribution in the RAW format. SMID: MID is a static video dataset consisting of frames captured with short exposures and corresponding ground truth obtained using long exposures. SID: The collection approach for SID is similar to that of SMID but introduces additional challenges with extreme situations. For both SMID and SID, we use full images and convert RAW data to RGB since our work focuses on LLE in the RGB domain. DAVIS2017: The utilization of the DAVIS dataset for LLE was initially proposed by [35]. It synthesizes low-light and normal-light video pairs with dynamic motions. In comparison to [35], we further incorporate the degradation of noise into low-light videos, in addition to invisibility, aligning more closely with real-world low-light videos. All training and testing splits adhere to the guidelines specified in the original papers.
Methods SDSD-indoor SDSD-outdoor Davis2017 Params Runtime
PSNR SSIM PSNR SSIM PSNR SSIM (millions) (seconds)
DP3DF [22] 28.9 0.880 27.24 0.794 - - 28.86 0.01
Ours 30.95 0.897 27.89 0.804 - - 33.32 0.01
SDSD [2] 25.86 0.738 25.35 0.765 - - 4.45 -
Ours 27.92 0.759 26.88 0.785 - - 6.82 -
StableLLVE [18] - - - - 26.39 0.977 - 0.01
Ours - - - - 28.04 0.981 - 0.02
Table 1: Quantitative comparison in LLVE task on SDSD-indoor, SDSD-outdoor, and Davis2017. Our method performs better than the baseline consistently. The runtime is computed with a size of 512$\times$960.
Figure 4: Visual comparison on different datasets. From top to down, they are Retinexformer in SDSDindoor, Restormer in SMID and SDSDoutdoor, and SNR in LOLv2-real. Our method yields better results.
@@ 4.1 Experimental Settings
Compared methods. To validate the effectiveness of our method, we conduct comparisons with a set of SOTA methods for both LLIE and LLVE tasks. For video tasks, we evaluate against DP3DF [26], SDSD [19], and StableLLVE [35], following their official settings and objectives. In the case of LLIE tasks, we include three representative frameworks: Retinexformer [1], Restormer [33], and SNR [25].
@@
Implementation details. We implement our framework in PyTorch and conducted experiments on an NVIDIA A100 GPU. Our implementation is based on the released code of the baseline networks, and we ensure the use of the same training settings for both the baseline and our method. The code will be made publicly available immediately following publication.
4.2 Quantitative Evaluation
Quantitative comparison on LLVE tasks. The quantitative comparisons for LLVE baselines and their versions with our depth guidance are presented in Table  1. Specifically, in DP3DF, our method demonstrates improvements of 2.05 in PSNR and 0.017 in SSIM on the SDSD-indoor dataset. Notably, within the SDSD framework, our method achieves enhancements of 2.06 in PSNR and 1.53 in SSIM on the SDSD-indoor dataset. Furthermore, in the StableLLVE framework evaluated on the Davis2017 dataset, our results exhibit a significant increase of 1.65 in PSNR, showcasing substantial improvement, particularly in various dynamic scenes. Additionally, we provide metrics for ”params” and ”running time” for both baselines and our method in Table  1. The results illustrate that the improvements achieved by our method do not come at the severe expense of efficiency.
Quantitative comparison on LLIE tasks. For the evaluation of LLIE, we conduct experiment on a variety of datasets, showing the effectiveness of our strategy more comprehensively and convincingly. The results are displayed in Table 2, where we can see that almost all baselines are improved by involving our strategy without additional heavy computation cost. The “-” iterms are placed due to the limited computation resources. In particular, for Retinexformer, our method showcases improvements of 1.84 in PSNR and 0.021 in SSIM on the SDSD-outdoor dataset. Within the Restormer framework, our method achieves enhancements of 0.012 in SSIM on the SDSD-outdoor dataset. Additionally, in the SNR framework evaluated on the LOLv2-synthetic dataset, our results demonstrate a notable increase of 1.83 in PSNR and 0.021 in SSIM, indicating substantial improvements across different scenarios.
Qualitative evaluation. We also provide the visual comparisons for the baseline and ours on different datasets, as shown in Figs. 4 and 5. As we can see, our method’s results are closer to the ground truth with natural illumination and color, fewer noise and artifacts.
@@ 4.2 Quantitative Evaluation
Methods SDSD-indoor SDSD-outdoor SMID LOLv2-real LOLv2-synthetic SID Params Runtime
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM (millions) (seconds)
Retinexformer [1] 25.00 0.894 25.69 0.780 28.66 0.805 - - 24.20 0.926 21.59 0.56 - 0.05
Ours 25.12 0.903 27.53 0.801 28.72 0.810 - - 24.97 0.928 22.17 0.566 - 0.06
Restormer [33] 27.11 0.923 29.53 0.838 - - 19.49 0.852 - - 22.99 0.533 26.13 0.15
Ours 27.40 0.924 29.91 0.850 - - 19.95 0.853 - - 23.77 0.634 31.41 0.21
SNR [25] 26.15 0.914 28.41 0.824 28.21 0.801 19.98 0.828 23.35 0.925 22.69 0.616 39.12 0.01
Ours 26.86 0.924 29.34 0.849 29.01 0.813 20.88 0.849 25.18 0.942 26.53 0.678 41.49 0.03
Table 2: Quantitative comparison in video task on SDSD-indoor, SDSD-outdoor, SMID, LOLv2-real, LOLv2-synthetic and SID. Our method performs better than the baseline consistently. The runtime is computed with a size of 512$\times$960.
Figure 5: Visual comparison on different datasets. From top to down, they are DP3DF in SDSDindoor and SDSDoutdoor, StableLLVE in Davis2017. Our method yields better results.
4.3 Ablation Study
In this section, we perform various ablation studies on the SDSD datasets, encompassing both indoor and outdoor subsets, to assess the significance of different components in our framework. Specifically, we evaluate the key components in HDGFFM through three ablated cases: (i) The location where the geometric factor is added. (ii) The method of fusion between image features and depth-aware features. (iii) Different weights for depth loss. The details of each case are elaborated below.
  • “Ours with Decoder”. As indicated in Sec.3.1 and Eq.2, the fusion in our framework occurs in the encoder of the target network $\mathcal{G}$. This choice is made because both the image content features and depth-aware features are extracted from the low-light input image, fostering homogeneity and enhancing fusion effects. To investigate the impact of alternative fusion points, we aim to test whether fusing image content features and depth-aware features in the decoder of $\mathcal{G}$ yields comparable results.
  • “Ours w/o Corre.”. As highlighted in Eq.8, we employ a correlation computation method via Sigmoid before the fusion of image content features and depth-aware features. This approach is designed to further mitigate inhomogeneity between these features. In this ablation setting, we omit the correlation computation, i.e., remove the Sigmoid from Eq.8. Consequently, the weight $w_{l}$ can have unlimited score ranges in this scenario.
  • “Ours with Add”. In a straightforward fusion approach, the two features are directly added without considering correlation computations. This setting neglects the incorporation of correlation computations between the features.
  • “Ours with $\lambda\times 5$” and “Ours with $\lambda\div 5$. As specified in Eq. 12, there is a hyper-parameter that governs the contribution of image reconstruction loss and depth prediction loss, denoted as $\lambda$. We aim to analyze the effects of varying the value of $\lambda$ on the final results.
As depicted in Tables 3 and 4, it is evident that the results of all ablation settings are inferior to our original implementation in the comparison section. In comparing “Ours” with “Ours with Decoder”, we validate the impracticality of fusing restored image content features and depth-aware features since the former is close to the normal-light output while the latter is close to the low-light inputs. Furthermore, comparing “Ours” with “Ours w/o Corre.” and ”Ours with Add” underscores the significance of our proposed “Correlation-based Cross Attention for Fusion”. Lastly, the consistent and even improved results across different values for $\lambda$ highlight the robustness of the hyperparameter choice in our framework. This emphasizes that researchers can choose optimal $\lambda$ values for different datasets.

### v2

@@ III-C Correlation-based Cross Attention for Fusion
$\small A_{l}=Softmax(W_{q}(f_{g,l})\times W_{k}(f_{g,l})^{T}\times\tau),$ (5)
where $W_{q}(f_{g,l})$ and $W_{k}(f_{g,l})$ has the shape of $C\times(H\times W)$, $T$ is the transpose operation, $\tau$ is a temperature hyper-parameter, $\times$ is the matrix multiplication operation, and $A_{l}$ has the shape of $C\times C$. We utilize a channel-wise attention map to maintain low computational costs, though an attention map with spatial dimensions is also compatible with our framework. The output of the self-attention is obtained as
$\small f^{\prime}_{g,l}=A_{l}\times W_{v}(f_{g,l}),$ (6)
where $f^{\prime}_{g,l}$ has the shape of $C\times(H\times W)$.
@@
$f^{\prime\prime}_{g,l}=\hat{A}_{l}\times\hat{W}_{v}(\hat{f}_{d,l}),$
where $\hat{A}_{l}$ and $f^{\prime\prime}_{g,l}$ have the same shape as that of $A_{l}$ and $f^{\prime}_{g,l}$.
Note that we adopt channel-wise attention due to its lower computational complexity compared to spatial-wise attention. We give a comprehensive complexity analysis here. The attention computation process consists of four main components: feature projection, attention map computation between the query and key features, the softmax operation, and the multiplication between the attention map and the value feature, as outlined in Eq 5 and Eq 6 of the main paper. Notably, the computational complexity remains the same for both self-attention (Eq 5/6) and cross-attention (Eq 7). We assume the input feature map has a shape of $H\times W\times C$.
  • Feature projection in Eq 5: Our projection layer includes $W_{q}$, $W_{k}$ (or $\hat{W}_{k}$), and $W_{v}$ (or $\hat{W}_{v}$). These are obtained via convolutional layers: the first layer is a $1\times 1$ convolution that transforms the channel dimension from $C$ to $3C$, and the second is a depthwise convolution with a kernel size of $3\times 3$ and group size $3C$, maintaining the channel dimension at $3C$. The computational cost of the first layer is $H\times W\times C\times 3C$, and that of the second layer is $H\times W\times 3\times 3\times 3C$. Therefore, the total computational complexity is $\mathcal{O}((27C+3C^{2})HW)$.
  • Attention map computation in Eq 5: For channel-wise attention, the attention map has a shape of $C\times C$, obtained by multiplying two features of sizes $C\times(HW)$ and $(HW)\times C$. The resulting complexity is $\mathcal{O}(C^{2}HW)$. In contrast, spatial-wise attention produces an attention map of size $(HW)\times(HW)$, formed by multiplying features of sizes $(HW)\times C$ and $C\times(HW)$, leading to a computational complexity of $\mathcal{O}(CH^{2}W^{2})$.
  • $\times\tau$ in Eq 5: The attention map is scaled by a temperature factor $\tau$, which involves an element-wise multiplication. This operation introduces an additional computational complexity of $\mathcal{O}(C^{2})$ for channel-wise attention and $\mathcal{O}(H^{2}W^{2})$ for spatial-wise attention, respectively.
  • Softmax operation in Eq 5: The softmax operation is applied row-wise to the attention map and involves three main steps: computing the exponent of each element, summing these exponents across the row, and dividing each element by the sum. Consequently, the computational complexity is $\mathcal{O}(3C^{2})$ for channel-wise attention and $\mathcal{O}(3H^{2}W^{2})$ for spatial-wise attention.
  • Multiplication between the attention map and value part in Eq 6: After obtaining the attention map, it is multiplied with the value feature to produce the output. For channel-wise attention, this step has a computational complexity of $\mathcal{O}(C^{2}HW)$; for spatial-wise attention, the complexity is $\mathcal{O}(CH^{2}W^{2})$.
In summary, the computational complexity of our channel-wise attention mechanism is $\mathcal{O}\big{(}(27C+3C^{2})HW+C^{2}HW+C^{2}+3C^{2}+C^{2}HW\big{)}=\mathcal{O}\big{(}(27C+5C^{2})HW+4C^{2}\big{)}$, where the dominant term is $\mathcal{O}((27C+5C^{2})HW)$ under the assumption that $C\ll H$ and $C\ll W$. In contrast, the complexity of the spatial-wise attention mechanism is $\mathcal{O}\big{(}(27C+3C^{2})HW+CH^{2}W^{2}+H^{2}W^{2}+3H^{2}W^{2}+CH^{2}W^{2}\big{)}=\mathcal{O}\big{(}(27C+3C^{2})HW+(4+2C)H^{2}W^{2}\big{)}$, which is significantly higher than that of the channel-wise attention.
As both $f_{g,l}$ and $\hat{f}_{d,l}$ are derived from the encoder process of $I_{d}$ and exhibit homogeneity, as discussed in Sec. III-B, $f^{\prime}_{g,l}$ and $f^{\prime\prime}_{g,l}$ also exhibit small heterogeneity between them. To further minimize the disparity between $f^{\prime}_{g,l}$ and $f^{\prime\prime}_{g,l}$, we introduce a method to model the correlation between these two sets of features via a fusion network, and then utilize this correlation as a weighting factor for the final fusion process. The correlation is obtained as
$\small w_{l}=Sigmoid(\mathcal{O}(f^{\prime}_{g,l}\oplus f^{\prime\prime}_{g,l})),$ (8)
where $\oplus$ is the channel concatenation operation and $\mathcal{O}$ is fusion network. The fusion output can be written as
$\small\bar{f}_{g,l}=w_{l}\cdot f^{\prime\prime}_{g,l}+f^{\prime}_{g,l}+f_{g,l},$ (9)
@@
III-D Training Strategy
In Eq 3, two objectives are considered: the restoration loss for enhancement and the depth supervision loss to obtain accurate depth priors. The restoration loss, denoted as $\mathcal{L}_{g}$, can employ the same loss terms as those used for the target model $\mathcal{G}$. On the other hand, the depth supervision loss is implemented as the $L_{2}$ distance between the predicted depth values and the ground truth depth information, as
$\small\mathcal{L}_{d}=\|\hat{d}_{d}-d_{n}\|_{2}.$ (11)
@@ III-D Training Strategy
$\small\mathcal{L}=\mathcal{L}_{g}+\lambda\mathcal{L}_{d},$ (12)
@@
where $\lambda$ is the loss weight, which is robust across different target scenarios. The training procedure of our method is summarized in Algorithm 1.
Algorithm 1 The training procedure of our method.
Training dataset with pairs $\{(I_{d},I_{n})\}$, initialized target enhancement network $\mathcal{G}$ with parameters $\theta_{g}$ and depth estimator for low-light data $\mathcal{F}$ with parameters $\theta_{f}$, pre-trained depth estimation network $\mathcal{D}$, current training iteration $T$, maximum training iteration $T_{max}$,
while $T<T_{max}$ do
  Sample training batch $(I_{d},I_{n})$ from the dataset
  Use the frozen $\mathcal{D}$ to predict $d_{n}=\mathcal{D}(I_{n})$
  Conduct the depth prediction from $I_{d}$ to obtain $\hat{d}_{d}$ as shown in Eq 1, and obtain the feature of each encoder layer as $f_{d}=\{f_{d,l}\},l\in[1,L]$
  Extract the image features from $I_{d}$ with the encoder of $\mathcal{G}$, as $f_{g}=\{f_{g,l}\},l\in[1,L]$
  Forward $\{f_{d,l}\},l\in[1,L]$ to the depth-aware embedding module $\mathcal{E}_{l}$, to obtain $\{\hat{f}_{d,l}\},l\in[1,L]$, as shown in Eq 4
  Obtain the fusion results of $\{f_{g,l}\},l\in[1,L]$ and $\{\hat{f}_{d,l}\},l\in[1,L]$ via Eqs 5, 6, 7, 8, 9, and 10, as $\{F_{g,l}\},l\in[1,L]$
  Compute the output images with the fusion results and the decoder of $\mathcal{G}$, as $\hat{I}_{n}$
  Compute the restoration loss as $\mathcal{L}_{g}$ with $\hat{I}_{n}$
  Compute the depth supervision loss as $\mathcal{L}_{d}$ with $\hat{d}_{d}$ and $d_{n}$, as shown in Eq 11
  Get the final loss as shown in Eq 12, and update the network of $\mathcal{G}$ and $\mathcal{F}$ simultaneously
  $T=T+1$
end while
trained model $\mathcal{G}$ and $\mathcal{F}$
The applicability of our strategy for LLVE. While our method is evidently well-suited for LLIE, we emphasize that it is also applicable to the LLVE task, although it is not specifically designed for it. However, it is rational for LLVE because the pre-trained depth estimator $\mathcal{D}$ we use is typically trained on video data, enabling it to produce stable depth estimates without noticeable flickering in normal videos. For example, see the consistent depth results demonstrated in the DepthAnything project (https://depth-anything.github.io/ and https://depth-anything-v2.github.io/). Therefore, using such models to provide frame-wise depth supervision does not cause significant temporal inconsistency. This justifies our application of the method to LLVE videos, resulting in the expected improvement.
IV Experiment
IV-A Experimental Settings
Datasets. We assessed the proposed framework using various LLIE/LLVE datasets, which include SDSD-indoor [7], SDSD-outdoor [7], LOLv2 [23], SMID [1], SID [38], and DAVIS2017 [39].
  • SDSD: The videos in SDSD were captured in dynamic pairs using electromechanical equipment.
  • LOLv2: LOLv2 is subdivided into LOLv2-real and LOLv2-synthetic, and they are commonly utilized in LLIE.
  • SMID and SID: SMID is a static video dataset consisting of frames captured with short exposures and ground truth obtained using long exposures. The collection approach for SID is similar to that of SMID but introduces additional challenges with extreme situations. For both SMID and SID, we use full images and convert RAW data to RGB since our work focuses on LLE in the RGB domain.
  • DAVIS2017: The utilization of the DAVIS dataset for LLE was initially proposed by [9]. It synthesizes low-light and normal-light video pairs with dynamic motions. In comparison to [9], we further incorporate the degradation of noise into low-light videos, in addition to invisibility, aligning more closely with real-world low-light videos.
All training/testing splits adhere to the guidelines specified in the original papers.
TABLE I: Quantitative comparison in the LLVE task on SDSD and SMID. Note that the batch size of $SD$ in the SMID dataset is 12 that is smaller than the original paper due to limited computation resources. “T.C.” means the temporal consistency evaluation using [40]
Methods SDSD-indoor SDSD-outdoor SMID
PSNR SSIM LPIPS RMSE T.C. PSNR SSIM LPIPS RMSE T.C. PSNR SSIM LPIPS RMSE T.C.
DP3DF [8] 28.90 0.880 0.235 9.290 4.16e-03 27.24 0.794 0.336 12.303 2.26e-03 25.60 0.727 0.339 14.225 6.21e-04
Ours 30.95 0.897 0.231 7.280 4.48e-03 27.89 0.804 0.343 10.708 2.28e-03 26.35 0.743 0.345 13.677 4.56e-04
SDSD [7] 25.81 0.749 0.328 13.468 4.53e-03 25.17 0.763 0.244 14.421 2.86e-03 24.13 0.715 0.294 16.870 4.86e-03
Ours 27.53 0.766 0.311 10.970 4.31e-03 27.26 0.794 0.210 11.352 2.06e-03 26.04 0.746 0.238 13.788 1.78e-03
TABLE II: Quantitative comparison in the LLVE task on the DAVIS2017 dataset.
Methods DAVIS2017
PSNR SSIM LPIPS RMSE
StableLLVE [9] 26.39 0.977 0.283 61.589
Ours 28.04 0.981 0.284 59.011
TABLE III: The experimental results on different categories of data, involving DP3DF on SDSD-outdoor.
Methods PSNR SSIM LPIPS RMSE
Building in Baseline 24.68 0.743 0.011 14.88
Building in ours 24.76 0.751 0.009 14.76
car in Baseline 29.65 0.843 0.004 8.74
car in ours 29.83 0.846 0.003 8.57
floor in Baseline 24.58 0.744 0.010 15.06
floor in ours 24.94 0.757 0.009 14.44
plant in Baseline 21.78 0.622 0.002 20.96
plant in ours 25.58 0.718 0.007 13.47
Average in Baseline 25.17 0.738 0.007 14.91
Average in ours 26.28 0.768 0.007 12.81
TABLE IV: The evaluation of temporal consistency using subjective scores. The best temporal consistency is marked as 5, while the worst is marked as 1.
Methods SDSD-indoor SDSD-outdoor
DP3DF 3.48 2.23
Ours 3.67 2.34
SDSD 3.92 2.85
Ours 4.11 3.49
[…]
Methods Parameter FLOPs Time
DP3DF 28.86 4.65 13.54
ours 33.32 4.94 21.44
SDSD 4.30 0.34 69.70
ours 6.67 1.32 483.00
StableLLVE 4.32 0.32 48.20
ours 8.60 0.70 209.50
Restormer 26.13 1.06 261.20
ours 31.41 1.20 309.64
SNR 39.12 0.18 37.50
ours 41.49 0.38 108.10
Retinexformer 1.61 0.13 85.80
ours 15.65 0.93 184.90
HVI-CIDNET 1.97 0.01 117.40
ours 2.75 0.02 134.10
IAT 0.09 0.01 33.70
ours 0.16 0.07 87.40
@@ IV-A Experimental Settings
Compared methods. To validate the effectiveness of our method, we conduct comparisons with a set of SOTA methods for both LLIE and LLVE tasks. For video tasks, we evaluate against DP3DF [8], SDSD [7], and StableLLVE [9], following their official settings and objectives except for batch size. For LLIE tasks, we include two representative frameworks: Restormer [42] and SNR [4]. Compared to the baseline, the modification brought by applying our method includes the extraction of depth-aware image features and fusion layers during inference, as well as the additional training of the depth estimator during training. Moreover, KinD [43], URetinex-Net++ [44], LLFlow-L-SKF [12], Retinexmamba [45], Wakeup-Darkness [27], Zero-DCE [46], RUAS [47], SCI [48], NeRCO [3], and Clip-LIE [49] are included to show the SOTA performance that can be achieved by our method.
@@
Implementation details. We implement our framework in PyTorch and conducted experiments on an NVIDIA A100 GPU and an NVIDIA A40 GPU. Our implementation is based on the released code of the baseline networks, and we ensure the use of the same training settings for both the baseline and our method except for batch size.
TABLE VI: Quantitative comparison in LLIE task on SDSD-indoor, SDSD-outdoor, SMID, LOLv2-real, LOLv2-synthetic, and SID. Our mothod outperforms almost all baselines consistently. The dataset to train “SNR” is the original dynamic SDSD dataset, which is the same as other methods in this paper and different from [4] using its static version. Note that the batch size of Restormer is smaller than the original paper due to limited computation resources.
@@
Methods SDSD-indoor SDSD-outdoor SMID LOLv2-real LOLv2-synthetic SID
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Retinexformer [50] 25.00 0.894 25.69 0.780 28.66 0.805 19.60 0.770 24.20 0.926 21.59 0.560
Ours 25.12 0.903 27.53 0.801 28.72 0.810 20.13 0.770 24.97 0.928 22.17 0.566
Restormer [42] 27.11 0.923 29.53 0.838 28.68 0.810 19.49 0.852 25.74 0.943 22.99 0.533
Ours 27.40 0.924 29.91 0.850 29.28 0.818 19.95 0.853 26.13 0.947 23.77 0.634
SNR [4] 26.15 0.914 28.81 0.848 28.21 0.801 19.98 0.828 23.93 0.929 22.69 0.616
SNR+SKF [12] 23.32 0.876 27.16 0.812 27.68 0.793 21.65 0.809 17.70 0.785 20.32 0.563
Ours 26.86 0.924 29.38 0.858 29.01 0.813 20.88 0.849 25.53 0.940 26.53 0.678
TABLE VII: The quantitative comparisons with the representative low-light image enhancement baseline and its version enhanced by other priors.
Methods SDSD-indoor SDSD-outdoor SMID LOLv2-real LOLv2-synthetic SID
LPIPS RMSE LPIPS RMSE LPIPS RMSE LPIPS RMSE LPIPS RMSE LPIPS RMSE
SNR [4] 0.068 14.061 0.139 9.961 0.148 10.903 0.066 30.265 0.034 19.317 0.202 20.471
SNR+SKF [12] 0.084 18.039 0.135 13.643 0.160 11.415 0.101 22.957 0.104 39.688 0.287 27.176
Ours 0.063 13.028 0.135 9.429 0.142 10.138 0.069 29.788 0.027 16.398 0.154 13.630
TABLE VIII: The quantitative comparisons between the representative low-light image enhancement baseline and its version enhanced by our method.
Methods SDSD-indoor SDSD-outdoor SMID
PSNR SSIM LPIPS RMSE PSNR SSIM LPIPS RMSE PSNR SSIM LPIPS RMSE
HVI-CIDNET [51] 22.71 0.903 0.142 24.225 26.07 0.806 0.292 12.570 26.41 0.775 0.223 13.612
Ours 25.81 0.913 0.122 16.323 27.02 0.866 0.215 11.359 26.64 0.614 0.264 11.862
IAT [52] 15.28 0.694 0.221 44.740 15.01 0.692 0.348 63.240 25.07 0.754 0.171 15.010
Ours 17.58 0.729 0.190 39.540 16.83 0.670 0.159 42.150 25.34 0.758 0.163 14.610
IV-B Comparison
Quantitative comparison on LLVE tasks. The quantitative comparisons for LLVE baselines and their versions with our depth guidance are presented in Tables. I and II. Specifically, in DP3DF, our method demonstrates improvements of 2.05dB in PSNR and 0.017 in SSIM on the SDSD-indoor dataset. Notably, within the SDSD framework, our method achieves an enhancement of 1.72dB in PSNR and 0.017 in SSIM on the SDSD-indoor dataset. Furthermore, in the StableLLVE framework evaluated on the DAVIS2017 dataset, our results exhibit a significant increase of 1.65dB in PSNR, showcasing substantial improvement, particularly in various dynamic scenes. In addition, we conduct a classification of the dataset to evaluate the performance on different regions. Using the pretrained semantic classification model CLIP [53] and manual refinement, we divide our evaluation dataset into different categories, including floor, building, plants, and car. As shown in Table III, our approach has better performance on all categories.
Moreover, we further employ the evaluation metric for temporal consistency, using warping error [40]. The results are displayed in Tables. I and II, showing that the temporal consistency of the results enhanced by our method is improved for most cases. Moreover, we believe it’s important to include a subjective evaluation of temporal consistency. To this end, we conducted a user study, inviting more than 30 participants. The results, shown in Table. IV, reveal that participants gave higher scores for temporal consistency in our method.
Additionally, we provide metrics for “model parameters”, “FLOPs”, and “running time” for both baselines and our method in Table. V. The results illustrate that the improvements achieved by our method do not come at the severe expense of efficiency.
Quantitative comparison on LLIE tasks. For the evaluation of LLIE, we conduct experiment on a variety of datasets, showing the effectiveness of our strategy more comprehensively and convincingly. The results are displayed in Table. VI, where we can see that almost all baselines are improved by involving our strategy without additional heavy computation cost. Within the Restormer framework, our method achieves enhancements of 0.012 in SSIM on the SDSD-outdoor dataset. Additionally, in the SNR framework evaluated on the LOLv2-synthetic dataset, our results demonstrate a notable increase of 1.6dB in PSNR and 0.011 in SSIM, indicating substantial improvements across different scenarios. Moreover, we include the comparisons with more baselines and metrics (LPIPS and RMSE), as displayed in Tables. VII and VIII. The baselines in Table VIII include HVI-CIDNET [51] and IAT [52].
Moreover, we note that inserting our strategy into HVI-CIDNet results in a decrease of 16.1% in SSIM and 4.1% in LPIPS on SMID. We explain this phenomenon as a result of using the same hyper-parameter $\lambda$ in Eq 12 across different datasets, showing the robustness of our method across diverse scenes. However, this hyper-parameter can be adjusted for different methods to achieve better results. For HVI-CIDNet, which includes an additional loss term in the HVS space alongside the RGB space, keeping $\lambda$ unchanged may relatively overweight this term and diminish the impact of the depth supervision loss $\mathcal{L}_{d}$ in Eq 12, leading to poorer depth estimation on low-light data. By increasing $\lambda$, we could observe improved performance that surpasses the baseline in all metrics.
Furthermore, we compare the optimal results of our method with a lot of representative low-light enhancement methods, including KinD, URetinex-Net++, LLFlow-L-SKF, Retinexmamba, Wakeup-Darkness, Zero-DCE, RUAS, SCI, NeRCO, and Clip-LIE. These methods contain the strategies of both supervised and unsupervised. The results are exhibited in Table IX, which demonstrate the SOTA performance brought by our strategy.
TABLE IX: The comparison between our method and several representative low-light image enhancement baselines.